sin150-cos150

sin150-cos150

Grade:12

1 Answers

Nikhil Upadhyay
8 years ago
sin 150=sin (180-30)=sin30=1/2
cos 150=cos(180-30)=-cos 30=-1/2
therefore, (1/2)-(-1/2)=(1/2)+(1/2)=1ANS
PLZ APPROVE.